OpenStudy (anonymous):

Which two numbers represent the value of Point W on the number line? 1 3/4 , 1.75 1 3/4 , 1.8 1 3/4 , 1.6 1 3/4 , 1.8 PLEASE HELP!!

OpenStudy (anonymous):

this is the number line |----|----|----|----|----|--W--|

OpenStudy (anonymous):

|----|---|---|---|--W| 0 1 2

OpenStudy (anonymous):

PLEASE HELP!!

OpenStudy (anonymous):

1 + 5/6 = 11/6 = 1.83 which i think is rounded down as one of your options to 1.8

OpenStudy (anonymous):

can any one help me?

OpenStudy (anonymous):

PLEASE HELP!!